Problem 12-02 AFN equation

Broussard Skateboard's sales are expected to increase by 15% from $7.0 million in 2016 to $8.05 million in 2017. Its assets totaled $4 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 7%, and the forecasted payout ratio is 60%. What would be the additional funds needed? Do not round intermediate calculations. Round your answer to the nearest dollar.$__________

Assume that an otherwise identical firm had $5 million in total assets at the end of 2016. Broussard's capital intensity ratio (A0*/S0) is _______ ( higher than / lower than / equal to ) than the otherwise identical firm; therefore, Broussard is ________ ( less / more / the same ) capital intensive - it would require __________ (a smaller / a larger / the same ) increase in total assets to support the increase in sales.
Answer #1

The standard formula for AFN is as shown below:

AFN = Additional Fund Needed

Ao So Lo So

Where, A0 = Total assets last year = $ 4,000,000

S0 = total sales last year = $ 7,000,000 mn

DeltaS = Change in sales = Expected sales - Last year's sales = 8.05 - 7 = $ 1.05 mn = 1,050,000

L0 = spontaneous liabilities = accounts payables + accruals = 450,000 + 450,000 = 900,000

PM = profit margin = 7%

S1 = Expected sales this year = 8,050,000

D = dividend payout ratio = 60%

Hence, AFN

)× 1,050,000-( 9,00,000 7,000, 000 4,000, 000 ( )× 1,050,000-0.07×8,050,000× 7,000, 000 (1- 0.60)

= $ 239,600

Capital intensity is measured by A0 / S0

For the same sales, quantum of asset deployed by Broussard = $ 4mn < $ 5 mn deployed by the identical firm. Hence,

Broussard's capital intensity ratio (A0*/S0) is lower than the otherwise identical firm; therefore, Broussard is less capital intensive - it would require a smaller increase in total assets to support the increase in sales.